solve the system by addition method
5x=4y-8
3x+7y=14

Answer by mananth(16946)

5 x + -4 y = -8 .............1
3 x + 7 y = 14 .............2

multiply (1)by 7
Multiply (2) by 4
35 x -28 y = -56
12 x + 28 y = 56
Add the two equations
47 x = 0
/ 47
x = 0.
plug value of x in (1)
5 x + -4 y = -8
0 + -4 y = -8
-4 y = -8
-4 y = -8
y = 2